# Topologically protected negative entanglement

### v3

@@ Abstract
The entanglement entropy encodes fundamental characteristics of quantum many-body systems, and is particularly subtle in non-Hermitian settings where eigenstates generically become non-orthogonal. In this work, we find that negative biorthogonal entanglement generically arises from topologically-protected non-orthogonal edge states in free fermion systems, especially within topological flat bands. Departing from previous literature which associated negative entanglement with exceptional gapless points, we show that robustly negative entanglement can still occur in gapped systems. Gapless 2D topological flat bands, however, exhibits novel $S_{A}\sim-\frac{1}{2}L_{y}^{2}\log L$ super volume-law entanglement behavior which scales quadratically with the transverse dimension $L_{y}$, independent of system parameters. This dramatically negative scaling can be traced to a new mechanism known as non-Hermitian critical skin compression (nHCSC), where topological and skin localization in one direction produces a hierarchy of extensively many probability non-conserving entanglement eigenstates across a cut in another direction. Our discovery sheds light on new avenues where topology interplays with criticality and non-Hermitian localization, unrelated to traditional notions of topological entanglement entropy. This topologically protected negative entanglement also manifests in the second Rényi entropy, which can be measured through SWAP operator expecation values.
@@
Introduction.– The entanglement entropy plays a crucial role in unveiling fundamental insights into the locality of quantum information. For instance, by scaling either according to the volume or area [1, 2, 3, 4], the entanglement entropy reveals whether quantum correlations pervade the entire system or remain localized. Intriguingly, numerous studies [5, 6, 7] have suggested that the presence of topological order can also be encoded in the entanglement entropy, as revealed by the presence of an additional constant term [8, 9] or discontinuities [10] in the scaling relation.
@@
In this work, extending into the non-Hermitian realm [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], we uncover a new manner in which topology can substantially influence entanglement entropy behavior. Specifically, we show that certain topological boundary states can exert a strongly non-local influence on the dominant entanglement behavior of the entire system, such probability non-conservation results in negative free-fermion entanglement entropy. A primary feature of non-Hermiticity is that the eigenstates of the Hamiltonian $H$ are generically no longer orthogonal, such that a biorthogonal basis with left and right eigenstates i.e. $H=\sum E_{n}\ket{\psi_{n}^{R}}\bra{\psi_{n}^{L}}$ | with $\bra{\psi^{L}_{m}}\psi^{R}_{n}\rangle=\delta_{mn}$ is needed to maintain orthogonality and retain the probabilistic interpretation of quantum mechanics [54, 55, 56, 57, 58, 59, 60, 61, 62]. Within this biorthogonal framework, recent studies have revealed that both bipartite entanglement entropy and Rényi entropy can manifest unexpected negative values [63, 64, 65, 66, 67, 68, 69, 70], attributable to the presence of geometric defectiveness at exceptional points (EPs) [66, 67, 68, 69, 70].
The first discovery in this work is that the presence of an EP [65] is not strictly a prerequisite for observing negative entanglement entropy values – instead, substantial non-orthogonality among the right eigenstates suffices, and spectacularly so when the non-orthogonality is enforced by a topological flatband. We investigate two 2D topological non-Hermitian models where the topological edge states in these models demonstrate nearly perfect overlap, while the overlap among bulk states remains minimal. Remarkably, this enables topological nontrivialness to be strategically employed to switch the negative entanglement entropy on or off.
@@
Most notably, we discover that non-Hermitian topological flatbands can give rise to unconventional negative super-volume-law entanglement scaling $S_{A}\sim-\frac{1}{2}L_{y}^{2}\log L$, where $L$ and $L_{y}$ are respectively the system dimensions normal and parallel to the entanglement cut. This enigmatic $-L_{y}^{2}$ scaling dependence arises not just due to the enhanced non-orthogonality of the states due to flatness of the band, but also the extensively many probability-nonconserving entanglement eigenstates that emerge due to the band criticality – in a new mechanism that we dub non-Hermitian critical skin compression (nHCSC).
@@
Negative entanglement from eigenstate non-orthogonality.– In the non-Hermitian context, the density operator that preserves its role as a probabilistic weight is the biorthogonal density matrix $\rho=\ket{\Psi^{R}}\bra{\Psi^{L}}$ |, where
@@
Crucially, this $\bar{P}$ operator crucially contain complete information of the $n$th-order Rényi entropy for free fermions  [^1] An analogous expression also holds for free bosons: $S_{A,boson}^{(n)}=\frac{1}{n-1}\text{Tr}\Big{[}\log\Big{(}\bar{P}^{n}-(\bar{P}-I)^{n}\Big{)}\Big{]}$.
$S_{A}^{(n)}=\frac{\log\text{Tr}(\rho_{A}^{n})}{1-n}$ $=$ $\frac{1}{1-n}\text{Tr}\Big{[}\log\Big{(}\bar{P}^{n}+(I-\bar{P})^{n}\Big{)}\Big{]}$ (3)
$=$ $\frac{1}{1-n}\sum_{i}\log(p_{i}^{n}+(1-p_{i})^{n}),\quad$
@@
$S_{A}$ $=$ $-\text{Tr}\rho_{A}\log\rho_{A}=-\text{Tr}[\bar{P}\log\bar{P}+(I-\bar{P})\log(I-\bar{P})]$ (4)
@@
where $I$ is the identity matrix and $p_{i}$ are the eigenvalues of $\bar{P}$. Physically, $p_{i}$ represents occupation probabilities subject to the restriction to subregion $A$, and are indeed real and within $[0,1]$ for Hermitian $H$. But in non-Hermitian settings, due to the non-conservation of probability across the subregion boundary, we can potentially have $p_{i}\gg 1$, as we shall show shortly. Importantly, $p_{i}>1$ contributes negatively to both $S_{A}^{(n>1)}$ and $S_{A}=\lim_{n\rightarrow 1}S_{A}^{(n)}$ entropies, as elaborated in the Supplement [73].
@@
where we have used $\Gamma_{A}^{2}=\Gamma_{A}$, we deduce that at least one of the $p_{i}$ must also have become very large, since the divergent elements in $U^{-1}$ do not in general cancel off with the small elements in $\Gamma_{A}U$ except in the case of vanishing entanglement cut $\Gamma_{A}=I$. However, we stress that even when $U$ is still full-rank with non-defective eigenspace, $\eta$ can already be extremely close to unity and contribute to negative entanglement.
Figure 1: Negative entanglement in the 4-band exceptional topological crossing model (Eq. (7)) under $y$-OBCs with $L_{y}=3$. (a) In the topologically non-trivial but Hermitian case ($\alpha=0,M=1.2,\delta=0$), the squared overlap $\eta(k)$ [Eq. (5)] of the topological eigenstates (blue) vanishes rigorously. (b) In the topologically trivial (gapped), non-Hermitian case ($\alpha=0.5\pi,M=3,\delta=2$), $\eta(k)$ of the closest bulk states (light blue) still vanishes essentially. (c) For the non-trivial Chern case ($\alpha=0,M=3,\delta=2$), perfect overlap i.e. $\eta(k)=1$ is reached where topological edge modes (red) cross. (d) The free fermion entanglement entropy $S_{A}$ (considering only the real part) for cases (a,b) respectively increases and saturates with system circumference $L$ as expected, but that from the topological exceptional crossing (c) exhibits a new $-\frac{1}{3}\log L$ scaling. The entanglement subregion is taken to be the half-cylinder with width $L/2$.
@@
Negative entanglement from exceptional topological crossing.– In this section and the next, we showcase two illustrative 2D systems where topologically protected edge states with $\eta\approx 1$ contribute negative entanglement to different degrees of success. We consider a twisted  [^3] Such that the discrete momentum takes values of $k=\pi/L,3\pi/L,...$, avoiding the singular $k=0$ point exactly. cylindrical geometry such that the $y$ direction contains $L_{y}$ unit cells, and is bounded so as to host topological edge states. The other translation-invariant (circumferential) direction contains $L$ unit cells, and we shall assign half i.e. $L/2$ of it to be the entanglement subregion $A$. $P(k)$ projects onto the occupied lower half bands with $\text{Re}[E(k)]<E_{F}=0$ unless otherwise stated. To emphasize the cut direction, we label its unit cells and momentum as $L$ and $k$ instead of $L_{x}$ and $k_{x}$.

### v4

# Topologically protected non-Hermitian super-volume-law entanglement

@@ Abstract
The entanglement entropy encodes fundamental characteristics of quantum many-body systems, and is particularly subtle in non-Hermitian settings where eigenstates generically become non-orthogonal. In this work, we find that negative biorthogonal entanglement generically arises from topologically protected non-orthogonal edge states in free fermion systems, especially within topological flat bands. Departing from previous literature which associated negative entanglement with exceptional gapless points, we show that robustly negative entanglement can still occur in gapped systems. Gapless 2D topological flat bands, however, exhibits novel $S_{A}\sim-\frac{1}{2}L_{y}^{2}\log L$ super-volume-law entanglement behavior which scales quadratically with the transverse dimension $L_{y}$, independent of system parameters. This dramatically negative scaling can be traced to a new mechanism known as non-Hermitian critical skin compression (nHCSC), where topological and skin localization in one direction produces a hierarchy of extensively many probability non-conserving entanglement eigenstates across a cut in another direction. Our discovery sheds light on new avenues where topology interplays with criticality and non-Hermitian localization, unrelated to traditional notions of topological entanglement entropy. This topologically protected negative entanglement also manifests in the second Rényi entropy, which can be measured through SWAP operator expectation values.
@@
The entanglement entropy plays a crucial role in unveiling fundamental insights into the locality of quantum information. For instance, by scaling either according to the volume or area [1, 2, 3], the entanglement entropy reveals whether quantum correlations pervade the entire system or remain localized. Intriguingly, numerous studies [4, 5, 6] have suggested that the presence of topological order can also be encoded in the entanglement entropy, as revealed by the presence of an additional constant term [7, 8] or discontinuities in the scaling relation.
@@
In this work, we extend the study of entanglement entropy into the non-Hermitian regime [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], where a primary feature is that the eigenstates of the Hamiltonian $H$ are generically non-orthogonal. To maintain orthogonality and preserve probabilistic interpretation of quantum mechanics, we employ a biorthogonal basis of left and right eigenstates–i.e., $H=\sum E_{n}\ket{\psi_{n}^{R}}\bra{\psi_{n}^{L}}$ | with $\bra{\psi^{L}_{m}}\psi^{R}_{n}\rangle=\delta_{mn}$ [29, 30, 31, 32, 33]. Within this biorthogonal framework, recent studies have revealed that both bipartite entanglement entropy and Rényi entropy can manifest unexpected negative values [34, 35, 36, 37, 38], attributable to the presence of geometric defectiveness at exceptional points (EPs) [35, 36, 37, 38]. Building upon these insights, we uncover a new mechanism by which topology in non-Hermitian systems substantially influence entanglement entropy behavior. Specifically, we show that certain topological boundary states can exert a strongly non-local influence on the dominant entanglement behavior of the entire system, such probability non-conservation leading to topologically protected negative free-fermion entanglement entropy.
@@
Most notably, in our non-Hermitian model featuring topological flat bands, we uncover an unconventional negative super-volume-law entanglement scaling, given by $S_{A}\sim-\frac{1}{2}L_{y}^{2}\log L$, where $L$ and $L_{y}$ are the system dimensions normal and parallel to the entanglement cut, respectively. This beyond-volume-law scaling reflects a novel quantum correlation structure that is fundamentally distinct from those observed in area-law and volume-law systems, and has rarely been reported in previous studies [39]. The enigmatic $-L_{y}^{2}$ scaling dependence arises not just due to the enhanced non-orthogonality of the states due to flatness of the band, but also the extensively many probability-nonconserving entanglement eigenstates that emerge due to the band criticality – in a new mechanism that we dub non-Hermitian critical skin compression (nHCSC).
Another key discovery of this work is that the presence of an EP is not strictly a prerequisite for observing negative entanglement entropy values – instead, substantial non-orthogonality among the right eigenstates suffices, and spectacularly so when the non-orthogonality is enforced by the topological flatband. We investigate two 2D topological non-Hermitian models where the topological edge states in these models demonstrate nearly perfect overlap, while the overlap among bulk states remains minimal. Remarkably, this enables topological nontrivialness to be strategically employed to switch the negative entanglement entropy on or off.
I Results
I.1 Negative entanglement from eigenstate non-orthogonality
@@
In the non-Hermitian context, the density operator that preserves its role as a probabilistic weight is the biorthogonal density matrix $\rho=\ket{\Psi^{R}}\bra{\Psi^{L}}$ |, where
@@
Crucially, this $\bar{P}$ operator contains complete information about the $n$th-order Rényi entropy for free fermions  [^1] An analogous expression also holds for free bosons: $S_{A,boson}^{(n)}=\frac{1}{n-1}\text{Tr}\Big{[}\log\Big{(}\bar{P}^{n}-(\bar{P}-I)^{n}\Big{)}\Big{]}$.
$S_{A}^{(n)}=\frac{\log\text{Tr}(\rho_{A}^{n})}{1-n}$ $=$ $\frac{1}{1-n}\text{Tr}\left[\log\left(\bar{P}^{n}+(I-\bar{P})^{n}\right)\right]$ (3)
$=$ $\frac{1}{1-n}\sum_{i}\log\left(p_{i}^{n}+(1-p_{i})^{n}\right),\quad$
@@
$S_{A}$ $=$ $-\text{Tr}\rho_{A}\log\rho_{A}=-\text{Tr}\left[\bar{P}\log\bar{P}+(I-\bar{P})\log(I-\bar{P})\right]$ (4)
@@
where $I$ is the identity matrix and $p_{i}$ are the eigenvalues of $\bar{P}$. Physically, each $p_{i}$ represents an occupation probability restricted to subregion $A$, and is real and bounded within $[0,1]$ for Hermitian Hamiltonian. In non-Hermitian settings, however, $p_{i}$ can become complex and take values beyond $[0,1]$, with magnitudes $|p_{i}|\gg 1$. This occurs due to the non-conservation of probability across the subregion boundary, as we shall demonstrate. Substituting such large $|p_{i}|$ into Eqs (3) and (4) can result in an unexpected negative entanglement entropy. In general, both Rényi entropies $S_{A}^{(n>1)}$ and von Neumann entropy $S_{A}$ calculated from complex $p_{i}$ can be complex. An important exception arises for PT-symmetric Hamiltonians, for which $S_{A}^{(n)}$ remains real (see Methods, subsection C.2 for details). The topological flat band model [Eq (7)] discussed later, serves as a paradigmatic example. For more general cases, such as the exceptional crossing model in Eq (20), where $\text{Im}(S_{A}^{(n)})\neq 0$, we show both the real and imaginary parts of the entropies, even though the the real part can be more directly measured.
@@
where we have used $\Gamma_{A}^{2}=\Gamma_{A}$, we deduce that at least one of the absolute values of $p_{i}$ must also have become very large, since the divergent elements in $U^{-1}$ do not in general cancel off with the small elements in $\Gamma_{A}U$ except in the case of vanishing entanglement cut $\Gamma_{A}=I$. However, we stress that even when $U$ is still full-rank with non-defective eigenspace, $\eta$ can already be extremely close to unity and contribute to negative entanglement.
Figure 1: Robustly negative entanglement entropy from the gapped topological flat bands of our 2-component Hamiltonian (Eq (7)), for parameters $B=1,t=0.8,a_{0}=1$ (with $b_{0}=1.2\neq 1$ to open up the gap). (a) For small cylinder length $L_{y}=5$ and $y$-direction OBCs, the energy spectrum $E(k)$ of the topological edge states |$\ket{\psi_{m1}},\ket{\psi_{m2}}$ (bolded) exhibits a small but visible gap, but their overlap factor $\eta(k)$ already approaches unity. (b) Upon increasing $L_{y}$ to 25, a topological flat band with exponentially small gap is observed within the non-trivial regime prescribed by Eq (8), with $\eta(k)\approx 1$ extremely closely. (c) With $y$-PBCs, the midgap flat band disappears and $\eta(k)$ deviates markedly from unity, even though the gap is still small. (d) For $y$-OBCs but not $y$-PBCs, the overlap $\eta(k)$ saturates very close to unity once $L_{y}\sim 10^{1}$. (e) The entanglement entropy scaling behavior $S_{A}$ for different $L_{y}$. Notably, as $L_{y}$ increases, $S_{A}$ decreases with $\log L$ more rapidly as $S_{A}\sim-(\alpha L_{y}+\beta)\log L$, with $\alpha\approx 0.6633$, $\beta\approx-4.1817$ according to obtained from numerical fitting (black). It also saturates at $S_{min}\sim-L_{y}$ when $L\gtrapprox L_{y}$.
I.2 Super-volume-law entanglement scaling induced by topological flat bands
@@
In this section, we showcase a representative 2D system where topological edge states in flat bands (with $\eta\approx 1$) give rise to negative entanglement entropy that scales with a super-volume-law. We consider a twisted  [^3] Such that the discrete momentum takes values of $k=\pi/L,3\pi/L,...$, avoiding the singular $k=0$ point exactly. cylindrical geometry such that the $y$ direction contains $L_{y}$ unit cells, and is bounded so as to host topological edge states. The other translation-invariant (circumferential) direction contains $L$ unit cells, and we shall assign half i.e. $L/2$ of it to be the entanglement subregion $A$. $P(k)$ projects onto the occupied lower half bands with $\text{Re}[E(k)]<E_{F}=0$ unless otherwise stated. To emphasize the cut direction, we label its unit cells and momentum as $L$ and $k$ instead of $L_{x}$ and $k_{x}$. The key idea is that the non-Hermitian skin effect (NHSE) localizes all states towards a common boundary, such that the states would exhibit extremely high overlap if they are furthermore macroscopically energetically degenerate, as in the topological flat band. From our results below, this holds true even if the flat bands are not strictly gapless or defective.
To realize topologically protected flat bands across an extended range of $k$, it suffices to consider a minimal 2-component 4442 topological bands are sufficient, since the NHSE pushes both of them onto the same boundary and makes them overlap significantly. Hamiltonian
with $k$-dependent asymmetric off-diagonal hoppings, where $a_{0},b_{0}$ and hopping distance $B>0$ are all real. This structure guarantees that the entanglement entropy $S_{A}$ remains real (see Methods, subsection C.2). Under OBCs in the $y$-direction, we treat the $x$-momentum $k$ as a parameter, such that Eq (7) effectively describes a 1D non-Hermitian SSH model, previously well studied using the generalized Brillouin zone (GBZ) method [9, 15]. Therefore, we know that the NHSE results in boundary state localization with a skin depth of $-2/\log[(b_{0}-\cos k)^{B}/a_{0}]$ and dynamically induces a real energy spectrum from non-Bloch PT-symmetry [45, 46, 47, 48]. Moreover, by computing the topological winding number defined via integration over the GBZ, we find almost-flat topological bands existing whenever
$|a_{0}(b_{0}-\cos k)^{B}|\leq t^{2},$ (8)
[…]
$\det[H_{\text{y-OBC}}(k)]=[a_{0}(b_{0}-\cos k)^{B}]^{L_{y}},$ (9)
and consequently, the energies $E_{e_{1}}(k)$ and $E_{e_{2}}(k)$ of the two topological edge modes satisfy:
$E_{e_{1}}(k)E_{e_{2}}(k)\approx\left[\frac{a_{0}(b_{0}-\cos k)^{B}}{t^{2}}\right]^{L_{y}}.$ (10)
This implies that the topological gap also exhibits an inverse exponential scaling, $\Delta=2|E_{e_{1}}|\sim(\text{Const.})^{-BL_{y}}$, which vanishes as $L_{y}\rightarrow\infty$, as derived in Methods subsection B.1, even though it never exactly closes. Accordingly, the band flatness depends exponentially on $BL_{y}$. Perfect gap closure is only possible for $|b_{0}|\leq 1$, and below we investigate these two cases separately.
Figure 2: Very robustly super-volume-law negative entanglement entropy $S_{A}$ from gapless topological flat bands of our 2-component Hamiltonian (Eq (7)) with parameters $b_{0}=1,t=0.5,a_{0}=2$. (a) Even for small $L_{y}=3,B=2$, a topological gapless flat band (bolded) with dispersion $E_{e}(k)\sim k^{BL_{y}}$ and overlap $\eta(k)\approx 1$ emerges around $k=0$ under $y$-OBCs. (b) As $k$ approaches 00, the skin effect ($r(k)\rightarrow 0$) strongly compresses the topological eigenstate onto site $1A$, leading to nearly perfect localization. (c) For this gapless case, occupancy eigenvalues $p_{i}$ (Eq (12)) of $\bar{P}$ dramatically exceed the $[0,1]$ interval due to non-Hermitian critical skin compression (nHCSC), with $p_{1},p_{2},p_{3},...$ (red, green, blue…) exhibiting a hierarchy of power-law dependencies with $L$. (d) The negative entanglement scaling is accurately approximated by $S_{A}\approx-\frac{1}{2}(BL_{y})^{2}\log L$ (Eq (13), black) across different $B,L_{y}$ combinations for sufficiently large cylinder circumference $L$. This is manifestly super-volume-law behavior with respect to the cylinder length $L_{y}$. (e) The coefficient of $\log L$ in the numerical $S_{A}$, extracted through the gradient of the $dS_{A}/d(\log L)$ plots (shades of blue), agrees well with $-\frac{1}{2}(BL_{y})^{2}$ (Eq (13)) when $L\gtrapprox 10^{2}$. At smaller $L$, super-volume-law dependence $\propto(BL_{y})^{2}$ still holds for smaller $BL_{y}$, albeit with a smaller coefficient.
1. Gapped topological flat bands ($b_{0}>1$). – Even though the topological bands are gapped, they become almost flat and touching as $L_{y}$ is increased, as depicted in Figs. 1(a,b). At large $L_{y}$, their gap becomes exponentially small within the topologically non-trivial region given by Eq (10), where the overlap factor $\eta(k)\approx 1$, indicating that |$\ket{\psi_{m1}}$ and |$\ket{\psi_{m2}}$ are nearly parallel. This suggests that states within an extensive continuum of $k$ closely approximate EPs, which is unexpected since the Hamiltonian $H(k,k_{y})$ in Eq (7) does not inherently feature EP crossings. As a comparison, for periodic boundary conditions(PBCs) in the $y$ direction [Fig. 1(c)], the topological flat band is absent, and the overlap $\eta(k)$ does not approach $1$ even as $L_{y}$ increased to a large value of $25$ where the (bulk) band gap becomes quite narrow. For $y$-OBCs, even at very small $L_{y}\approx 5$ number of layers, $\eta(k)$ is already very close to one [Fig. 1(d)]; at larger $L_{y}$, $\eta(k)$ converges exponentially to 1 despite the system being physically gapped.
This strong topological flat band-induced non-orthogonality ($\eta\approx 1$) is manifested in a strongly negative bipartite entanglement entropy $S_{A}$. As shown in Fig. 1(e), $S_{A}$ scales negatively with the cylinder circumference $\log L$, with a gradient that grows with its length $L_{y}$. From Methods, subsection B.3, the exact dependence is established as $S_{A}\sim-(\alpha L_{y}+\beta)\log L$, where $\alpha\approx 0.6633$, $\beta\approx-4.1817$ as obtained from numerical fitting. Notably, this linear dependence on $L_{y}$ does not arise trivially because the length of the entanglement cut scales with $L_{y}$, since it is contributed only by the topological modes whose number do not scale extensively with system length. Rather, it arises because the band flatness scales exponentially with $L_{y}$. That said, for a given $L_{y}$, the entanglement entropy $S_{A}$ saturates at negative lower bound (Methods, subsection B.3) $S_{min}\sim-L_{y}\log[a_{0}(b_{0}-1)^{-B}]$ because the system is ultimately gapped, such that the overlap $\eta$ at $k_{0}=\pi/L$ (nearest point to $k=0$) does not approach arbitrarily close to 1 with increasing $L$. For PBCs, the bulk gap also results in the saturation of $S_{A}$ at a positive value, as depicted by the starred grey trend in Fig. 1(e).
2. Gapless topological flat bands ($b_{0}=1$) with super-volume-law negative entanglement. – Finally, we discuss the most intriguing case where $\det[H_{\text{y-OBC}}(k)]=0$ at $k=0$ [Eq (9)], such that the topological band gap vanishes exactly [Fig. 2(a)]. Even though its band structure looks superficially similar to the $b_{0}>1$ case with exponentially small gap [Fig. 1], its entanglement entropy exhibits a surprising super-volume-law dependence $S_{A}\sim-\frac{1}{2}B^{2}L_{y}^{2}\log L$, proportional not to the cylinder length $L_{y}$, but to the square of it. While the first power of $L_{y}$ can be attributed to the exponentially high topological band flatness as before, the additional second power of $L_{y}$ emerges from an uniquely new 2D phenomenon which we call non-Hermitian critical skin compression (nHCSC).
To understand the nHCSC, recall that all states are pushed to the cylinder’s edges due to the $y$-hopping asymmetry in our model Eq (7). In particular, the left and right biorthogonal topological eigenstates exhibit exponential spatial-$y$ profiles $\bra{\psi_{edge}^{L}(k,y)}\sim r(k)^{-y}$ and |$\ket{\psi_{edge}^{R}(k,y)}\sim r(k)^{y}$ with $r(k)=\sqrt{|(1-\cos k)^{B}/a_{0}|}<1$, and are localized at edges $y=L_{y}$ and $y=1$, respectively. Since these are the two eigenstates that exhibit the highest overlap $\eta(k)$, they dominate the entanglement contribution in $\bar{P}$. Due to their criticality (gaplessness), $r(0)=0$ and the state is perfectly skin-localized (see Fig. 2(b)) for the $k=0$ topological mode in the $x$-direction along the cylinder edge.
Ordinarily, this perfect edge localization (or “compression”) only leads to irreversible 1D non-Bloch dynamics [49] and singular generalized Brillouin zones  [50, 51, 52]. However, in our 2D topological entanglement context, it also causes the occupied band projector
to diverge for matrix blocks $y<y^{\prime}$, with strongest divergence in $[P(k)]_{1,L_{y}}\sim r(k)^{-L_{y}}\sim(1-\cos k)^{-BL_{y}/2}$. Notably, the most strongly divergent contribution $\sim k^{-BL_{y}}$ from $(1-\cos k)^{-BL_{y}/2}$ does not dominate the total negative entanglement; of also substantial significance are the entire set of divergent terms $k^{-BL_{y}+2},k^{-BL_{y}+4},k^{-BL_{y}+6},...$ from the sub-leading terms in the expansion of $(1-\cos k)^{-BL_{y}/2}$, as well as other $[P(k)]_{y,y^{\prime}}$ (see Methods, subsection B.2). Consequently, distinct from ordinary EP crossings [34, 35], essentially the entire set of $\bar{P}$ eigenvalues $p_{i}$, diverges with $L$:
$p_{i}\approx\text{Re}(p_{i})\sim\{L^{\frac{BL_{y}-1}{2}},L^{\frac{BL_{y}-1}{2}-1},L^{\frac{BL_{y}-1}{2}-2},...\},$ (12)
where the imaginary parts Im$(p_{i})$ remain negligible, as shown in Fig. 2(c). This is the main consequence of nHCSC, which hinges on both the edge compression of the eigenstates and its criticality (vanishing of $r(k)$). The hierarchy of these divergent occupancy eigenvalues is shown in Fig. 2(c): upon closer inspection, subdominant $p_{2},p_{3}\sim L^{\frac{BL_{y}-1}{2}-1},L^{\frac{BL_{y}-1}{2}-2}$ eigenvalues (green, blue) are observed in addition to the dominant $p_{1}\sim L^{\frac{BL_{y}-1}{2}}$. Summing over them, the total entanglement entropy scales like (Methods, B.2)
$S_{A}$ $=$ $-\sum_{p_{i}}p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i})$ (13)
$\approx$ $-\frac{1}{2}(BL_{y})^{2}\log L.$
This strongly negative entanglement $S_{A}$ is plotted in Fig. 2(d) for various $B,L_{y}$, and can be as low as $-70$ for reasonably large $L_{y}=4,B=2$. By examining the slope of $S_{A}$ with respect to the universal $\log L$ factor, it is numerically confirmed in Fig. 2(e) that the super-volume-law quadratic coefficient $-(BL_{y})^{2}/2$ accurately holds for across a wide range of $BL_{y}$ as long as $L\gtrapprox 10^{2}$ (even though moderately large $L\sim\mathcal{O}(10)$ suffices when $BL_{y}$ is also of $\mathcal{O}(10)$).
I.3 Prospects for measuring negative entanglement through the second Rényi entropy
As established in subsection C.1 of the Methods, the negative entanglement also manifests generically as negative Rényi entropy. Below, we outline a scheme for measuring the second Rényi entropy, defined in the biorthogonal basis as:
$S_{A}^{(2)}=-\log\text{Tr}\big{[}(\rho_{A})^{2}\big{]},$ (14)
A known approach [53, 54, 55] for measuring the second Rényi entropy or quantum purity involves the SWAP operator, which exchanges two copies of a quantum state:
[…]
A commonly used corollary [56] is $\text{Tr}(\text{SWAP}\rho_{1}\otimes\rho_{2})=\text{Tr}(\rho_{1}\rho_{2})$, from which the second Rényi entropy can be calculated from the expectation value of the SWAP operator on two-copies of the many body state  [57] as:
[…]
$=$ $\text{Tr}(\rho_{A}^{2})\quad=e^{-S_{A}^{(2)}},$
[…]
$\big{|}\bra{\Psi^{L}}\otimes\bra{\Psi^{L}}\text{SWAP}_{A}\ket{\Psi^{R}}\otimes\ket{\Psi^{R}}\big{|}$ | $=$ $e^{-\text{Re}[S_{A}^{(2)}]}.$ (17)
Therefore, to measure the (real part of the) biorthogonal second Rényi entropy, $\text{Re}[S_{A}^{(2)}]$, the most mathematically direct way would be to prepare two copies of the ground state of the Hamiltonian $H$ as |$\ket{\Psi^{R}}\otimes\ket{\Psi^{R}}$, applying the SWAP operator in subregion $A$, and then measuring its overlap with the ground state |$\ket{\Psi^{L}}\otimes\ket{\Psi^{L}}$ of another Hamiltonian $H^{\dagger}$. This approach could potentially be implemented using programmable quantum computers [58, 59, 60, 61]. Post-selection, which has been used in measuring negative conditional entropy [62], will also be useful in simulating the non-Hermiticity [63, 64, 25].
[…]
or
[…]
Thus, to measure $\big{|}\bra{\Psi^{L}}\otimes\bra{\Psi^{L}}\text{SWAP}_{A}\ket{\Psi^{R}}\otimes\ket{\Psi^{R}}\big{|}$ | as given in Eq (17), one feasible strategy is to prepare a superposition state with known amplitudes $c_{1},c_{2}$ in a physical system:
[…]
and then measure the expectation value $\bra{\Psi}\text{SWAP}_{A}\ket{\Psi}$. The left-right overlap terms (i.e. Eq (17)) can subsequently be calculated by subtracting the contributions from $\langle\text{SWAP}_{A}\rangle_{RR}$ and $\langle\text{SWAP}_{A}\rangle_{LL}$, which can also be independently measured. Moreover, the OBC spectrum of our model, as specified in Eq. (9), is purely real, leading to identical energies for |$\ket{\Psi^{R}}$ and |$\ket{\Psi^{L}}$ and thus facilitating the preparation of their superposition.
Beyond the approach described above, other potentially feasible ways for observing negative entanglement can involve directly measuring the reduced density matrix to calculate the entanglement entropy through quantum state tomography  [65, 66, 67], as well as measuring the non-local correlations in phononic crystals [68], from which the entanglement entropy can be inferred. Related quantum simulations through quantum Monte Carlo approaches [55, 69] and ultracold atomic optical lattices [70, 71, 53, 54, 72] can also reveal the associated quantum correlations.
I.4 Discussion
We have discovered a super-volume-law scaling of free fermion entanglement entropy in non-Hermitian, topologically non-trivial systems–particularly those featuring topological flat bands. Specifically, the entropy scales quadratically as $S_{A}\sim-\frac{1}{2}L_{y}^{2}\log L$, despite the entanglement cut being only of length $L_{y}$. This highly unconventional scaling goes beyond the well known area and volume law, and to our knowledge, has no analogue in Hermitian systems. We identify the origin of this behavior as a new mechanism we call non-Hermitian critical skin compression (nHCSC), where the criticality of highly degenerate NHSE-compressed topological modes gives rise to an extensive hierarchy of probability non-conserving $\bar{P}$ eigenstates that gives rise to even stronger negative $S_{A}$. Moreover, we find that that this negative scaling feature relies on substantial eigenstate overlap around the Fermi surface, which is less stringent than the previously suggested requirement of an exceptional crossing [34]. As such, macroscopically degenerate flatbands resulting from simultaneous topological and non-Hermitian skin localization can lead to strongly negative entanglement entropy, even in the presence of a small gap. Importantly, this negative entanglement also applies to the Rényi entropy which can be physically measured as suggested in previous subsection C, placing topology as a potentially practical control knob for probability non-conserving negative entanglement.
Figure 3: Negative entanglement in the 4-band exceptional topological crossing model (Eq (20)) under $y$-OBCs with $L_{y}=3$. (a) In the topologically non-trivial but Hermitian case ($\alpha=0,M=1.2,\delta=0$), the squared overlap $\eta(k)$ [Eq (5)] of the topological eigenstates (blue) vanishes rigorously. (b) In the topologically trivial (line-gapped along Re$(E)=0$), non-Hermitian case ($\alpha=0.5\pi,M=3,\delta=2$), $\eta(k)$ of the closest bulk states (light blue) still vanishes essentially. (c) For the non-trivial Chern case ($\alpha=0,M=3,\delta=2$), perfect overlap i.e. $\eta(k)=1$ is reached where topological edge modes (red) cross. (d) The free fermion entanglement entropy $S_{A}$ (considering only the real part) for cases (a,b) respectively increases and saturates with system circumference $L$ as expected, but that from the topological exceptional crossing (c) exhibits a new $-\frac{1}{3}\log L$ scaling. The entanglement subregion is taken to be the half-cylinder with width $L/2$.
II Methods
II.1 A comparative 4-band model with exceptional topological crossing
II.1.1 Negative entanglement from exceptional topological crossing